HELP i have two triangles ans the order from the bottom left corner is A the point is B and the right corner C inbetween the B AND C line there is the number 18 and between A AND C there is an x the other one is a triangle same order but D E AND F between E AND F There is a 12 and between D AND F there is a 6 . . WHAT DO THEY WANT?? the directions say triangle ABC~ triangle DEF and solve for x
Answer by venugopalramana(3286)

HELP i have two triangles ans the order from the bottom left corner is A the point is B and the right corner C inbetween the B AND C line there is the number 18 and between A AND C there is an x the other one is a triangle same order but D E AND F between E AND F There is a 12 and between D AND F there is a 6 . . WHAT DO THEY WANT?? the directions say triangle ABC~ triangle DEF and solve for x
THIS MEANS BC=18......AC=X.........EF=12.......DF=6
TRIANGLE ABC IS SIMILAR TO TRIANGLE DEF...
THEN AB/DE=BC/EF=CA/FD..HENCE..
18/12=X/6
12X=6*18
X=18*6/12=9